# A quantum information theoretic analysis of reinforcement learning-assisted quantum architecture search

### v1

@@ Abstract
In the field of quantum computing, variational quantum algorithms (VQAs) represent a pivotal category of quantum solutions across a broad spectrum of applications. These algorithms demonstrate significant potential for realising quantum computational advantage. A fundamental aspect of VQAs involves formulating expressive and efficient quantum circuits (namely ansatz) and automating the search of such ansatz is known as quantum architecture search (QAS). RL-QAS involves optimising QAS using reinforcement learning techniques. This study investigates RL-QAS for crafting ansatzes tailored to the variational quantum state diagonalization problem. Our investigation includes a comprehensive analysis of various dimensions, such as the entanglement thresholds of the resultant states, the impact of initial conditions on the performance of RL-agent, the phase change behavior of correlation in concurrence bounds, and the discrete contributions of qubits in deducing eigenvalues through conditional entropy metrics. We leverage these insights to devise an optimal, admissible QAS to diagonalize random quantum states. Furthermore, the methodologies presented herein offer a generalised framework for constructing reward functions within RL-QAS applicable to variational quantum algorithms.
@@ I Introduction
QAS typically consists of two parts. First, a template of the circuits is built, called the ansatz. The ansatz can have parametric quantum gates, e.g., rotation angles. Then, these parameters are determined via the variational principle using a classical optimiser in a feedback loop. Algorithms constructed via this technique are called variational quantum algorithms (VQA) McClean et al. (2016); Cerezo et al. (2021). QAS can also used to determine non-parametric circuits as an approach for quantum program synthesis Sarkar (2022). In this research, however, we will focus on QAS in the context of VQA.
@@
The ansatz design is critical, as it directly influences the expressivity and efficiency of the quantum solution. Given the vast potential configuration space of quantum circuits, the process of identifying a suitable ansatz is a challenging problem. In previous works such as Ostaszewski et al. (2021); Kundu et al. (2024); Patel et al. (2024); Du et al. (2022), the search for the quantum circuits that solve various VQA has been automated using reinforcement learning (RL) techniques Sutton and Barto (2018). We particularly call quantum ansatzes capable of expressing the solution of VQA an admissible ansatz. The automated search for such admissible ansatz using the RL-agent is termed reinforcement learning-assisted quantum architecture search (RL-QAS). In this setting, after a certain number of episodes, the RL agent returns possible configurations of quantum ansatz that solve the problem. For $E_{tot}$ episodes, the agent returns $E_{s}$ successful episodes, and in $E_{tot}-E_{s}$ episodes, the agent fails to find a configuration within a specific predefined accuracy and depth. It has been shown in Ostaszewski et al. (2021); Wu et al. (2023); Kundu et al. (2024); Patel et al. (2024); Kundu (2024) that classical machine learning-driven QAS algorithms outperform the state-of-art structures of ansatz and effectively optimise the number of parameters, depth, and impact of noise in VQAs.
@@
The use-cases of such admissible ansatzes cover applications in quantum information theory LaRose et al. (2019); Cerezo et al. (2021); Tan and Volkoff (2021); Cerezo et al. (2020); Kundu and Miszczak (2022), chemistry Kandala et al. (2017); Mustafa et al. (2022); Delgado et al. (2021) and combinatorial optimization Khairy et al. (2020); Liu et al. (2022); Glos et al. (2022). Yet, a thorough investigation of such admissible ansatzes is still missing in recent literature. It has been previously shown Patel et al. (2024) that in RL-QAS methods, it is crucial to feed proper information about the problem to the agent to enhance the agent’s performance. One of the ways of providing this information is by engineering a reward function. In RL, the formulation of the reward function and its engineering can be sparse and dense. The choice of optimal reward function depends on trial and error by investigating the performance of the agent under various reward configurations.
@@
In this article, we investigate the admissible ansatzes proposed by the RL-agent from the perspective of quantum information theory Wilde (2013). For the investigation, we specifically focus on the reinforcement learning enhanced variational quantum state diagonalisation (RL-VQSD) problem Kundu et al. (2024). In the VQSD algorithm, the problem-inspired ansatz is based on density matrix exponentiation Lloyd et al. (2014), but its depth and number of gates increase exponentially with the number of qubits. Hence, searching for an optimal ansatz (with a minimum number of gates and depth) that diagonalises a quantum state is an open research problem. In RL-VQSD, the search for the optimal configuration of ansatz is done by using a double deep-Q network (DDQN). See Appendix B for further details on the VQSD algorithm, the agent and environment specifications. The RL-agent proposes $E_{s}<E_{tot}$ admissible ansatzes that solves the VQSD problem.
@@
Our investigation shows that the concurrence of the $s$ ansatzes lies between an upper and lower bound, independent of the different initialisation of the double deep-Q network (DDQN), where each initialisation of DDQN corresponds to a different random initialisation of the weights. Moreover, we show that the upper and lower bounds are anti-correlated with respect to the increment in the initial state entanglement. The anti-correlation turns into a mild, and later strong, correlation when the initial state entanglement goes beyond $0.322$, indicating a phase change in correlation between the upper and lower bounds of the concurrence. The phase transition provides us with an in-depth insight into the relation between the input and the output of the VQSD algorithm. Further, our investigation reveals that the RL-agent can generate high concurrence admissible ansatz with fewer 2-qubit gates and circuit depth compared to ansatzes with very low concurrence. Furthermore, we quantify the contribution of individual qubits of the RL-ansatz using conditional quantum entropy. Using such a measure, we observe a correlation between entanglement and quantum entropy, revealing why RL-VQSD can efficiently find the largest eigenvalues but fails to find the smallest ones.
@@ II Contributions
    Investigating admissible ansatzes of reinforcement learning based quantum architecture search problem from a quantum information theory perspective. We focus on the RL-VQSD algorithm, where the proposed ansatzes diagonalise random quantum states.
@@
    Demonstrating that the entanglement of the states generated by the RL-agent-proposed admissible ansatzes lies within a lower and upper concurrence bound.
@@
    The upper and lower bounds of concurrence remain consistent across various weight initialisation of the DDQN.
@@
    Bounds are anti-correlated with respect to increasing entanglement until concurrence surpasses $0.322$; after this point, the anti-correlation turns into correlation. Hence, we introduce the concurrence $0.322$ as the phase change point.
  5. 5.
    Investigation of ansatzes in upper and lower bounds, showing optimal ansatz for VQSD generates high concurrence states.
  6. 6.
    RL-agent requires fewer gates and circuit depth to produce highly entangled admissible ansatzes than weakly entangled ansatzes. We utilise this information to minimise RL-agent and environment (which is the configuration of ansatz) interaction, thereby optimising the convergence time of the RL-VQSD.
  7. 7.
    We evaluate individual qubits’ contribution in diagonalising 2-qubit quantum states using conditional quantum entropy.
  8. 8.
    Our investigation shows that the qubits have equal contributions on average in diagonalising random quantum states.
  9. 9.
    We further investigate the correlation in conditional entropy between qubits for different inferred eigenvalues. It requires mild anti-correlation for the first two largest eigenvalues and strong anti-correlation for the smallest eigenvalues.
  10. 10.
    The majority of admissible ansatzes lie in a mild anti-correlation regime, explaining VQSD’s limitation in finding the smallest eigenvalues.
@@ III.1.1 DDQN weight independent entanglement bound
To benchmark the performance of the RL agent for a specific problem, it is ideal to take multiple initialisation of the DDQN (see Appendix B for a brief description).
@@
Table 1: In this table, we present the average and standard deviation of the concurrence of the final state after evolution through the circuit for five different initialisations of the DDQN wrights for different quantum seed states.
To efficiently investigate the upper and lower bounds of concurrence Hill and Wootters (1997); Wootters (1998) produced by the RL ansatz, we first sample different random quantum states, then diagonalize them using RL-VQSD Kundu et al. (2024) for five different initialisation of DDQN weights. Thereafter, the agent suggests different possible configurations for the RL-ansatz, which diagonalizes the sampled random quantum states. For a specific state, we collect such valid quantum circuits and find the concurrence of the state obtained after evolving through the RL-ansatz. It should be noted that finding the concurrence of the evolved state indicates how much the RL-ansatz induces entanglement to the input sampled state. See Appendix C for further analysis of the sampled random quantum states.
@@ III.1.2 Anti-correlation between entanglement bounds
Figure 1: Reciprocal behaviour in the upper and lower bounds of concurrence with respect to the increasing concurrence in the input random state. A clear anti-correlation between entanglement minima and maxima which breaks down after $0.3$. From the interpolating function, the Pearson correlation coefficient (PCC) for the first eleven points is -0.982, while that for the remaining points is 0.049. At this point, the correlation after the first eleven points changes from -0.055 to 0.049. There is a phase transition from weak anti-correlation to weak correlation.
Corresponding to a predefined threshold, the agent proposes different architectures of quantum circuits that diagonalise the quantum state. To observe the amount of entanglement generated by the RL-ansatz we calculate the concurrence of the state after passing through the RL-ansatz, just before the dephasing operation. By sorting out the quantum circuit with maximum and minimum entanglement, we observe their variation with the entanglement of the input state. It has been proved in the previous section that the upper and lower bounds of entanglement are independent of different initialisation of DDQNs; hence, we for further benhamarking we can constrain ourselves to just one initialisation of the DDQN weights.
In Figure 1, we illustrate the variation of the entanglement bounds with the entanglement of the input quantum state. To investigate the dependency, we use the Pearson correlation coefficient (PCC) Lee Rodgers and Nicewander (1988) as a quantifier of correlation. See Appendix A for a brief description. We see that the upper and lower bounds of entanglement are initially strongly anti-correlation when the input state entanglement is below $0.3$ and then strongly correlate afterwards, indicating a phase change in the upper and lower bounds of entanglement.
Figure 2: The PCC indicating Pearson correlation coefficient is calculated over an interval $i<k<j$. The coefficient is described as in Equation 1, where $i$ is minimum, $j$ is maximum concurrence, and $k$ splits the intermediate values of concurrence into two parts. Hence, $i$ and $j$ are fixed. The $w_{ik}\in[0,1]$ is the weight of the first interval which defines how much each of the two intervals contributes when varying $k$. We vary the cardinality of the two intervals, causing variation in the weights. In the figure, we consider $i=0.119$ and $j=0.402$ and when $k=0.322$, i.e. $i<x\leq 0.322$, the weight corresponds to the first interval.
@@
$\eta^{k}_{w_{ik}}=w_{ik}\textrm{PCC}^{ik}_{ul}+(1-w_{ik})\textrm{PCC}^{kj}_{ul},$ (1)
where $w_{ik}\in\{0,1\}$ and $\textrm{PCC}^{ik}_{ul}$, $\textrm{PCC}^{kj}_{ul}$ are the value of PCC between the upper and lower bounds of concurrence of the ansatz for the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$ respectively.

### v2

@@ Abstract
In the field of quantum computing, variational quantum algorithms (VQAs) represent a pivotal category of quantum solutions across a broad spectrum of applications. These algorithms demonstrate significant potential for realising quantum computational advantage. A fundamental aspect of VQAs involves formulating expressive and efficient quantum circuits (namely ansatz) and automating the search of such ansatz is known as quantum architecture search (QAS). RL-QAS involves optimising QAS using reinforcement learning techniques. This study investigates RL-QAS for crafting ansatzes tailored to the variational quantum state diagonalisation problem. Our investigation includes a comprehensive analysis of various dimensions, such as the entanglement thresholds of the resultant states, the impact of initial conditions on the performance of RL-agent, the phase change behaviour of correlation in concurrence bounds, and the discrete contributions of qubits in deducing eigenvalues through conditional entropy metrics. We leverage these insights to devise entanglement-guided admissible ansatz in QAS to diagonalise random quantum states using optimal resources. Furthermore, the methodologies presented herein offer a generalised framework for constructing reward functions within RL-QAS applicable to variational quantum algorithms.
@@ I Introduction
QAS typically consists of two parts. First, a template of the circuits is built, called the ansatz. The ansatz can have parametric quantum gates, e.g., rotation angles. Then, these parameters are determined via the variational principle using a classical optimiser in a feedback loop. Algorithms constructed via this technique are called variational quantum algorithms (VQA) McClean et al. (2016); Cerezo et al. (2021). QAS can also be used to determine non-parametric circuits as an approach for quantum program synthesis Sarkar (2022). In this research, however, we will focus on QAS in the context of VQA.
@@
The ansatz design is critical, as it directly influences the expressivity and efficiency of the quantum solution. Given the vast potential configuration space of quantum circuits, the process of identifying a suitable ansatz is a challenging problem. In Ostaszewski et al. (2021); Kundu et al. (2024); Patel et al. (2024); Du et al. (2022), the search for the quantum circuits that solve various VQA has been automated using reinforcement learning (RL) techniques Sutton and Barto (2018). We particularly call ansatzes capable of solving a VQA an admissible ansatz. The automated search for such admissible ansatz using the RL-agent is termed reinforcement learning-assisted quantum architecture search (RL-QAS). In this setting, after a certain number of episodes, the RL agent returns possible configurations of quantum ansatz that solve the problem. For $E_{tot}$ episodes, the agent returns $E_{s}$ successful episodes, and in $E_{tot}-E_{s}$ episodes, the agent fails to find a configuration within a specific predefined accuracy and depth. It has been shown in Ostaszewski et al. (2021); Wu et al. (2023); Kundu et al. (2024); Patel et al. (2024); Kundu (2024) that classical machine learning-driven QAS algorithms outperform the state-of-art structures of ansatz and effectively optimise the number of parameters, depth, and impact of noise in VQAs.
@@
The use-cases of such admissible ansatzes cover applications in quantum information theory LaRose et al. (2019); Cerezo et al. (2021); Tan and Volkoff (2021); Cerezo et al. (2020); Kundu and Miszczak (2022), chemistry Kandala et al. (2017); Mustafa et al. (2022); Delgado et al. (2021) and combinatorial optimization Khairy et al. (2020); Liu et al. (2022); Glos et al. (2022). Yet, a thorough investigation of such admissible ansatzes is still missing in recent literature. It has been previously shown Patel et al. (2024) that in RL-QAS methods, it is crucial to feed proper information about the problem to the agent to enhance the agent’s performance. One of the ways of providing this information is by engineering a reward function. In RL, the formulation of the reward function and its engineering can be sparse or dense. The choice of optimal reward function depends on trial and error by investigating the performance of the agent under various reward configurations.
@@
In this article, we investigate the admissible ansatzes proposed by the RL-agent from the perspective of quantum information theory Wilde (2013). For the investigation, we specifically focus on the reinforcement learning enhanced variational quantum state diagonalisation (RL-VQSD) problem Kundu et al. (2024). In the VQSD algorithm LaRose et al. (2019), the problem-inspired ansatz is based on density matrix exponentiation Lloyd et al. (2014), but its depth and number of gates increase exponentially with the number of qubits. Hence, searching for an optimal ansatz (with a minimum number of gates and depth) that diagonalises a quantum state is an open research problem. In RL-VQSD, the search for the optimal configuration of ansatz is done by using a double deep-Q network (DDQN). See Appendix B for further details on the VQSD algorithm, the agent and environment specifications. The RL-agent proposes $E_{s}<E_{tot}$ admissible ansatzes that solves the VQSD problem.
@@
Our investigation shows that the concurrence of the $E_{s}$ ansatz lies between an upper and lower bound, independent of the different initialisation of the double deep-Q network (DDQN), where each initialisation of DDQN corresponds to a different random initialisation of the weights. Moreover, we show that the upper and lower bounds are anti-correlated with respect to the increment in the initial state entanglement. The anti-correlation turns into a mild, and later strong, correlation when the initial state entanglement goes beyond $0.322$, indicating a phase change in correlation between the upper and lower bounds of the concurrence. The phase transition provides us with an in-depth insight into the relation between the input and the output of the VQSD algorithm. Further, our investigation reveals that the RL-agent can generate high concurrence admissible ansatz with fewer 2-qubit gates and circuit depth compared to ansatzes with very low concurrence. This is used as a guiding principle to design a two-part ansatz, where an entanglement enhancing initial block adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$. Furthermore, we quantify the contribution of individual qubits of the RL-ansatz using conditional quantum entropy. Using such a measure, we observe a correlation between entanglement and quantum entropy, revealing why RL-VQSD can efficiently find the largest eigenvalues but fails to find the smallest ones.
@@ II Contributions
    Demonstrating that the entanglement of the states generated by the RL-agent-proposed admissible ansatzes lies within a lower and upper concurrence bound. The upper and lower bounds of concurrence remain consistent across various weight initialisation of the DDQN. Furthermore, we show that these bounds are anti-correlated with respect to increasing entanglement until concurrence surpasses $0.322$; after this point, the anti-correlation turns into correlation. Hence, we introduce the concurrence $0.322$ as the phase change point.
@@
    Investigation of ansatzes in upper and lower bounds reveals that the optimal ansatz for VQSD generates strongly entangled states. The RL-agent requires fewer gates and circuit depth to generate such states to produce strongly entangled admissible ansatzes than weakly entangled ansatzes. We then utilise this information to minimise RL-agent and environment (which is the configuration of ansatz) interaction, thereby optimising the convergence time of the RL-VQSD. Specifically, we propose a two-part ansatz, where an entanglement enhancing initial block adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$.
@@
    We evaluate individual qubits’ contribution in diagonalising 2-qubit quantum states using conditional quantum entropy. Our investigation shows that the qubits have equal contributions on average in diagonalising random quantum states.
@@
    We further investigate the correlation in conditional entropy between qubits for different inferred eigenvalues. It requires mild anti-correlation for the first two largest eigenvalues and strong anti-correlation for the smallest eigenvalues. Most admissible ansatzes lie in a mild anti-correlation regime, explaining VQSD’s limitation in finding the smallest eigenvalues.
@@ III.1.1 DDQN weight independent entanglement bound
To benchmark the performance of the RL-agent for a specific problem, it is ideal to take multiple initialisations of the DDQN (see Appendix B for a brief description).
@@
Table 1: The average and standard deviation of the final state’s maximum (max. concurrence) and minimum concurrence (min. concurrence) after evolution through the circuit for five different initialisations of the DDQN weights for different quantum seed states. To obtain the results, we run RL-VQSD to diagonalize $2$-qubit random quantum states for $10000$ episodes. The agent then proposes $E_{s}<10000$ admissible ansatzes, which generate $E_{s}$ states after the random quantum state evolves them. The max. (min.) concurrence is attributed to the ansatz that generates a state with the highest (lowest) concurrence following the evolution of the random quantum state.
To efficiently investigate the upper and lower bounds of concurrence Hill and Wootters (1997); Wootters (1998) of the admissible ansatz, we calculate the concurrence of admissible ansatz, which is defined by the entanglement induced by the admissible ansatz when initialised with |$\ket{00}$. We first sample different random quantum states, then diagonalise them using RL-VQSD Kundu et al. (2024) for five different initialisations of DDQN weights. Thereafter, the agent suggests different possible configurations for the RL-ansatz, which diagonalises the sampled random quantum states. For a specific state, we collect such valid quantum circuits and find the concurrence of the state obtained after evolving through the RL-ansatz. It should be noted that finding the concurrence of the evolved state indicates how much the RL-ansatz induces entanglement to the input sampled state. See Appendix C for further analysis of the sampled random quantum states, where we show that in qiskit, random_density_matrix even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low.
@@ III.1.2 Anti-correlation between entanglement bounds
Corresponding to a predefined threshold, the agent proposes different architectures of quantum circuits that diagonalise the quantum state. To observe the amount of entanglement generated by the RL-ansatz we calculate the concurrence of the state after passing through the RL-ansatz, just before the dephasing operation. By sorting out the quantum circuit with maximum and minimum entanglement, we observe their variation with the entanglement of the input state. It has been proved in the previous section that the upper and lower bounds of entanglement are independent of different initialisation of DDQNs; hence, for further benchmarking, we can constrain ourselves to just one initialisation of the DDQN weights.
Figure 1: Reciprocal behaviour in the upper and lower bounds of concurrence with respect to the increasing concurrence in the input random state. For the first eleven points, the Pearson correlation coefficient (PCC) is $-0.982$, denoting a strong anti-correlation between the upper and lower bounds of concurrence. The strong anti-correlation turns into a mild correlation after concurrence of the initial state surpasses $0.322$. The correlation after the first eleven points changes from $-0.055$ to $0.049$. There is a phase transition from weak anti-correlation to weak correlation.
In Figure 1, we illustrate the variation of the entanglement bounds with the entanglement of the input quantum state. To investigate the dependency, we use the Pearson correlation coefficient (PCC) Lee Rodgers and Nicewander (1988) as a quantifier of correlation. See Appendix A for a brief description. We see that the upper and lower bounds of entanglement are initially strongly anti-correlation when the input state entanglement is below $0.322$ and then strongly correlate afterwards, indicating a phase change in the upper and lower bounds of entanglement.
Figure 2: We present the phase transition of correlation between the upper and lower bounds of concurrence for the admissible RL-VQSD ansatz. To show this, we split the total range of the concurrence of the input state in two intervals $\mathbbm{u}_{1}\coloneqq[i,k)$ and $\mathbbm{u}_{2}\coloneqq[k,j]$, where $k$ is a free parameter and plot the variation of the Pearson correlation coefficient between the upper and lower bound on the concurrence of the ansatz corresponding to these two intervals. We denote by $\textrm{PCC}_{ik}$ and $\textrm{PCC}_{kj}$ the correlation coefficients in the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$ respectively. The phase transition point is observed at $k=0.322$ where the concurrence bounds change from mild anti-correlation to mild correlation for the interval $\mathbbm{u}_{2}$ while remaining strongly anti-correlated in the interval $\mathbbm{u}_{1}$.
@@
$\eta^{k}_{w_{ik}}=w_{ik}\textrm{PCC}_{ik}+(1-w_{ik})\textrm{PCC}_{kj},$ (1)
where $w_{ik}\in\{0,1\}$ and $\textrm{PCC}_{ik}$, $\textrm{PCC}_{kj}$ are the value of PCC between the upper and lower bounds of concurrence of the ansatz for the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$ respectively. For the sake of simplicity we denote $\textrm{PCC}_{ik}$ as $\textrm{PCC}_{ik}^{ul}$ and also $\textrm{PCC}_{kj}$ as $\textrm{PCC}_{kj}^{ul}$.
@@
We observe that depending on the values of $w_{ik}$, $\eta^{k}_{w_{ik}}$ can take values either $\textrm{PCC}_{ik}$ or $\textrm{PCC}_{kj}$. For $\eta^{k}_{w_{ik}}=\textrm{PCC}_{ik}$, we obtain the amount of (anti-)correlation between the lower and upper bounds of concurrence of the ansatz for the interval $\mathbbm{u}_{1}$. While for $\eta^{k}_{w_{ik}}=\textrm{PCC}_{kj}$, we obtain the amount of (anti-)correlation for the interval $\mathbbm{u}_{2}$. We leave proving $\eta^{k}_{w_{ik}}$ is a valid measure of correlation for arbitrary values of $w_{ik}\in(0,1)$ as an open problem.
We present in Figure 2, the variation of $\eta_{0}^{k}$ (in blue) and $\eta_{1}^{k}$ (in red) for different values of $k$. We observe that $\eta_{1}^{k}\approx-1~{}\forall k\in(i,j)$. This indicates strong cumulative anti-correlation between the upper and lower bounds of concurrence. For $\eta_{0}^{k}$, we observe a gradual change of $\text{PCC}_{kj}$ from -0.582 to 1. We note that $\eta_{0}^{k}$ changes from $-0.0451$ to $0.279$ for $k\in[0.313,0.379]$ with $\eta_{0}^{k}=0$ at $k=k_{\ast}=0.322$. We call this point the phase transition point of correlation between upper and lower bounds of concurrence for the VQSD ansatz. At values of $k<k_{\ast}$, there is anti-correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$. While for values of $k>k_{\ast}$, there is (anti-)correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $(\mathbbm{u}_{2})~{}\mathbbm{u}_{1}$.
III.1.3 Enhancing the performance of agent based on initial state entanglement
Figure 3: The RL-agent can generate an ansatz with strong entanglement with fewer 2-qubit gates and circuit depth compared to the ansatzes with weak entanglement. As the main aim of RL-VQSD is to diagonalise unitary with a small number of gates and depth, our analysis shows that if we particularly focus on admissible ansatzes near the upper bound of concurrence, we can not only reduce the computational cost of RL-VQSD severely but find optimal admissible ansatzes. The numbers in the bars correspond to the amount of entanglement generated by the RL-ansatz when initialised in the vacuum state.
Our investigation reveals that the admissible ansatzes corresponding to the upper bound of concurrence has, on average, a shorter depth of $13.23$ and a smaller number of gates of $18.5$ compared to the admissible ansatzes corresponding to the lower bound of concurrence where the depth is $13.9$ and number of gates is $19$ over $25$ random quantum states. Moreover, Figure 3 makes it more prominent that the minimum number of gates and the depth of the admissible ansatzes are smaller for the upper bound of concurrence than the lower. This observation is crucial in reducing the subspace of all possible admissible ansatzes. The main motivation of RL-QAS in VQAs is to find ansatzes that solve a VQA with a minimum number of gates and circuit depth; in the case of RL-VQSD, this observation can help us to narrow down the search for an optimal admissible ansatz in the two following ways.
a. Entanglement dependent initialisation of RL-state
In RL-VQSD Kundu et al. (2024), the RL-state that encodes the ansatz initiates from an empty circuit (i.e., in circuit 2 $U_{\text{RL-agent}}=\mathds{1}$). We showed that the optimal admissible ansatz, on average, lies in the upper bound of concurrence, motivating us to initialise the quantum state (that is diagonalised using RL-VQSD) closest to the upper bound of entanglement. To do this, the ansatz is divided into two parts